Common stock valuelong dashVariable growth   

Lawrence​ Industries' most recent annual dividend was ​$2.18 per share ​(D 0 = $ 2.18​), and the​ firm's required return is 13​%. Find the market value of​ Lawrence's shares when dividends are expected to grow at 10​% annually for 3​ years, followed by a 5​% constant annual growth rate in years 4 to infinity.
Answer #1

D1=(2.18*1.1)=2.398

D2=(2.398*1.1)=2.6378

D3=(2.6378*1.1)=2.90158

Value after year 3=(D3*Growth rate)/(Required return-Growth rate)

=(2.90158*1.05)/(0.13-0.05)

=38.0832375

Hence market value=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.398/1.13+2.6378/1.13^2+2.90158/1.13^3+38.0832375/1.13^3

=$32.59(Approx)